I am stuck on correct answer for:
if AX - B = 0 and A is not equal to 0, then x =

I am not using any specific book. I am just practicing..

Answer by funmath(2933)

First be careful not to change from upper case to lower case like you did when you posted your question. It didn't matter on this problem because there was not a mixture of upper and lower case letters. But some formulas do and "A" would not be the same as "a."
:
Second, we want "x" all by itself on one side of the equal sign. What ever we do to accomplish that, we have to do to both sides of the equal sign.
:
Ax-B=0 (B doesn't have an x and is - so add B to both sides.)
Ax-B+B=0+B
Ax+0=B
Ax=B (There's nothing between A and x, that means they're being multiplied, so we divide A on both sides.)
Ax/A=B/A
x=B/A
Happy Calculating!!!